Give the x-intercept and y-intercept for the equation 8x - 2y = 24. (Put in ordered pair form)
Found 2 solutions by nerdybill, actuary:
Answer by nerdybill(7384)

To find the x-intercept, set y=0 and solve for 'x':
8x - 2y = 24
8x - 2(0) = 24
8x - 0 = 24
8x = 24
x = 3
Therefore, the x-intercept is at (3,0)
.
To find the y-intercept, set x=0 and solve for 'y':
8x - 2y = 24
8(0) - 2y = 24
0 - 2y = 24
-2y = 24
y = -12
Therefore, the y-intercept is at (0,-12)

Answer by actuary(112)

The y-intercept occurs when x = 0 so solve the given equation for y when x = 0.
0*8-2*y = 24 Therefore, y = -12. In ordered pair form, {0,-12}
The x-intercept occurs when y = 0 so solve the given equation for y = 0.
8*x - 2*0 = 24 Therefore, 8*x = 24 or x = 3. In ordered pair form, {3,0}.